The boiling point of water is 100.0°C at 1 atmosphere.

How many grams of iron(II) iodide (309.7 g/mol), must be dissolved in 258.0 grams of water to raise the boiling point by 0.300°C? Refer to the table for the necessary boiling or freezing point constant.

Solvent Formula Kb (°C/m) Kf (°C/m)
Water H2O 0.512 1.86
Ethanol CH3CH2OH 1.22 1.99
Chloroform CHCl3 3.67
Benzene C6H6 2.53 5.12
Diethyl ether CH3CH2OCH2CH3 2.02


___?___ g iron(II) iodide.
Answer #1

I for FeI2 is 3 as it dissociates into 1 Fe2+ and 2 I-.

use:

Δ Tb = i*Kb*mb

0.3 = 3.0*0.512 *mb

mb= 0.1953 molal

mass, m = 2.58*10^2 g

= 0.258 kg

use:

molality = mass of solute/(molar mass of solute * mass of solvent in Kg)

0.1953 = massOfSolute/(309.7*0.258)

massOfSolute = 15.606 g

Answer: 15.61 g